OpenStudy (anonymous):

Given f(x) = –3x – 7, what is f(–4)? A. 40 B. 5 C. –14 D. –1

OpenStudy (anonymous):

f(x) = –3x – 7 (other example) f(-2) = –3(-2) – 7=6-7=-1 (other example) f(5) = –3(5) – 7=-15-7=-22 f(-4)=? (plug in 4 for x and solve, just like I showed on f(-3) and f(5) )

OpenStudy (anonymous):

I meant to say plug in -4 for x...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thank you for the help :D

OpenStudy (anonymous):

No problem! Good luck !